# Edge-Cloud Collaborative Computing on Distributed Intelligence and Model Optimization: A Survey

### v2

@@ II-B Distributed Intelligence Framework
Collaborative intelligence in edge-cloud systems leverages both edge and cloud resources to achieve optimal performance by distributing intelligence across the architecture. The distributed intelligence framework adheres to three key principles: task parallelism, data parallelism, and model parallelism. Task parallelism executes different parts of complex AI tasks concurrently on separate devices [45], represented as $T=\{t_{1},t_{2},...,t_{n}\}$, where $T$ represents the overall task and $t_{i}$ represents individual sub-tasks. Data parallelism trains identical models on different data subsets distributed across multiple devices [11], expressed as $D=\{d_{1},d_{2},...,d_{m}\}$, where $D$ represents the complete dataset and $d_{i}$ represents the data subset on device $i$. Model parallelism deploys different parts of large models on separate devices [46], formulated as $M=\{m_{1},m_{2},...,m_{p}\}$, where $M$ represents the complete model and $m_{i}$ represents the model partition on device $i$. Edge intelligence focuses on real-time processing with low latency [47], while cloud intelligence handles complex computations and large-scale training [48]. Model partitioning strategies balance workloads by considering model complexity, data size, and resource constraints [49]. Resource allocation algorithms dynamically assign resources based on real-time demands [50], while performance optimization techniques like model compression enhance efficiency [51].
@@ II-C Resource Types and Constraints
Computation, communication bandwidth, and storage capacity are fundamental resources in edge-cloud systems, each with inherent constraints impacting system design and performance [39]. Specifically, computation resources, denoted as $C_{k}$ for node $k$, represent processing power; communication bandwidth, $B_{ij}$, limits data transfer rates between nodes $i$ and $j$, directly affecting latency; and storage capacity, $S_{k}$, dictates the data volume node $k$ can store, influencing data-intensive applications [52]. Managing these constrained and heterogeneous resources across diverse devices, under dynamic workloads and with limited edge resources, especially energy for battery-powered devices [40], presents significant challenges.
@@ III-A Model Compression
Deploying sophisticated AI models on resource-constrained edge devices presents substantial challenges due to limited computational power, memory, and energy resources. Various model compression techniques [2, 56, 57] have emerged as critical solutions to reduce model size and computational demands while maintaining acceptable performance levels. Table II provides a comprehensive comparison of recent approaches across three primary compression categories: pruning, quantization, and knowledge distillation. Notably, pruning techniques demonstrate significant improvements in inference latency and energy efficiency by strategically eliminating less important parameters, as further detailed in Sec. III-A1. Meanwhile, quantization methods, explored in Sec. III-A2, reduce parameter precision requirements, with implementations achieving remarkable energy efficiency in healthcare applications. Knowledge distillation approaches, discussed in Sec. III-A3, enable smaller models to effectively mimic larger counterparts, with recent innovations specifically optimized for efficient LLM inference across edge-cloud boundaries. Across these diverse techniques, a common pattern emerges wherein most successful implementations leverage both edge and cloud resources collaboratively, highlighting the synergistic nature of modern compression strategies in distributed computing environments.
TABLE II: Summary of model compression techniques in AI-driven ECCC AIoT systems.
@@ III-A1 Pruning
Strategically eliminating redundant model parameters, pruning techniques reduce size and complexity in ECCC, enabling efficient AI model deployment on resource-constrained edge devices [59]. We explore various strategies targeting different architectural aspects, such as pruning entire neurons or filters based on performance contribution, as in the two-phase split computing framework (TSCF) [58], or focusing on individual weights. Applied during or after training, dynamic and static pruning balance accuracy, and computational cost. For instance, Hybrid SD [59] employs structural pruning of the Stable Diffusion Model’s U-Net to maintain image quality on edge devices, while the adaptive pruning-split federated learning (PSFed) method [60] optimizes semantic coder updates in a SemCom-assisted SEC framework, reducing training delay and energy consumption.
@@ III-A2 Quantization
Lower-precision representation of model parameters effectively reduces both memory footprint and computational cost [62], making quantized models particularly suitable for resource-constrained edge devices. Instead of 32-bit floating-point representations, quantization employs lower precision formats such as 8-bit integers, consequently diminishing memory requirements and accelerating computations on these lower-precision numbers. For example, [61] explored model quantization for edge implementation in arrhythmia monitoring to reduce energy consumption. Additionally, QuAsyncFL [63] integrates quantization with asynchronous federated learning to enhance communication efficiency in cloud-edge-terminal collaborations within the AIoT domain.
@@ III-A3 Knowledge Distillation
Model compression through knowledge distillation (KD) trains a smaller ”student” model to mimic a larger ”teacher” model’s behavior [68]. This process enables the student model to achieve comparable performance with reduced complexity, facilitating efficient deployment in resource-constrained edge-cloud environments [56]. Instead of learning from ground truth labels, the student model learns the teacher model’s output distribution, providing richer information than hard labels and enabling more nuanced representations and better generalization [65]. KD benefits ECCC by placing the teacher model in the resource-rich cloud and the distilled student model on edge devices [6], optimizing overall system performance. Agglomerative federated learning (FedAgg) [67] uses a bridge sample based online distillation protocol (BSBODP) for knowledge transfer between nodes across different tiers, enabling larger model training while maintaining privacy and flexibility. Guidance-based knowledge transfer (GKT) [68] employs a larger LLM as a ”teacher” to generate guidance prompts for a smaller ”student” model, enhancing accuracy and speed without fine-tuning. Growth-adaptive distillation [6] automatically adjusts the student model structure for efficient, low-power edge deployment. In cloud-edge collaborative reinforcement learning, algorithm-independent knowledge distillation [69] accelerates online RL agent convergence by leveraging pre-trained cloud models and offline agent interactions. A federated bi-directional knowledge distillation strategy [70] enhances fault detection in manufacturing by integrating quality knowledge from both cloud and edge, strengthening interaction between sub-processes and hierarchical levels.
@@ III-B Model Adaptation
Effectively adapting AI models to new tasks, datasets, or evolving data distributions is crucial within the dynamic landscape of ECCC systems. To achieve this adaptability, three primary learning paradigms are widely employed: transfer learning (TL), which accelerates learning on target tasks by leveraging knowledge from pre-trained models; federated learning (FL), which facilitates privacy-preserving, distributed training across multiple clients; and continual learning (CL), which enables models to learn sequentially from evolving data streams without catastrophic forgetting. The interplay and distinct roles of these paradigms in fostering adaptability within ECCC systems are conceptually depicted in Fig. 4. Furthermore, Table III presents a comparative summary of recent research efforts utilizing TL, FL, and CL, highlighting their specific approaches, application domains, and system characteristics within ECCC. Following this, the subsequent sections delve into the transfer learning in Sec. III-B1, federated learning in Sec. III-B2, and continual learning in Sec. III-B3.
@@
Method Year Venue Type Domain Edge Cloud ECC Model Comp.
@@
Note for ’Type’: TL = Transfer Learning, FL = Federated Learning, CL = Continual Learning
@@ III-B1 Transfer Learning
Within edge-cloud environments, adapting pre-trained models to new tasks or datasets often leverages transfer learning [72]. This approach reduces training time and data requirements by building upon existing knowledge, thereby improving deployment efficiency, especially in resource-constrained systems [65]. For example, in surface defect detection, a model pre-trained on a general image dataset can be fine-tuned with a smaller, specialized dataset to identify specific defects [72]. Similarly, in cloud-device collaborative learning for multimodal large language models (MLLMs), techniques like adapter-based knowledge distillation (AKD) facilitate knowledge transfer from larger cloud models to smaller, compressed device models [65], maintaining performance while minimizing computational burden. In cloud-edge collaborative fault diagnosis, transfer learning addresses data heterogeneity and distribution shifts by fine-tuning source models trained on diverse user data to bridge distribution gaps and enable effective diagnosis on target datasets [75].
III-B2 Federated Learning
As a pivotal technique for distributed model training, federated learning [89] preserves data privacy by enabling clients to train local models on their own datasets and aggregate updates (e.g., gradients) on a central server to refine a global model [86]. This collaborative approach, by keeping sensitive data localized, not only addresses privacy concerns but also enhances the global model’s generalization capabilities by leveraging diverse datasets [80]. In edge-cloud systems, FL efficiently and securely leverages distributed computational resources [82], as exemplified by frameworks like FedCAE [87] for machine fault diagnosis and QuAsyncFL [63] for AIoT applications. Specifically, SFETEC [95] optimizes FL for thing-edge-cloud environments by splitting model training to reduce communication overhead and address non-IID data. For enhanced security against poisoning attacks, FlexibleFL [88] evaluates uploaded parameters and participant contributions. Additionally, optimizing participant selection and learning topologies, as in [85], improves FL efficiency, while integrating FL with technologies like blockchain [83] further bolsters security and trustworthiness in collaborative learning.
@@ III-B3 Continual Learning
In ECCC, adapting models to new data or tasks incrementally is crucial for lifelong learning [96]. This adaptability enables models to evolve with dynamic data and emerging tasks, addressing the core challenge of catastrophic forgetting, where new information disrupts previously acquired knowledge [96]. Several strategies, including regularization, replay, and dynamic architectures, mitigate this issue, ensuring robust and adaptable systems. Specifically, continual learning enables digital twins to synchronize with their evolving physical counterparts and sensor data in edge computing, providing accurate predictions [92, 91]. For instance, in intelligent transportation systems (ITS), cloud-edge collaborative continual adaptation frameworks, leveraging visual prompts and knowledge distillation, enhance the generalization of lightweight edge models for object detection in dynamic environments [93]. Similarly, online management frameworks, such as EdgeC3, optimize model aggregation and dynamic offloading for continuous data streams in edge-cloud collaborative continual learning [94]. Cross-edge federated continual learning (Cross-FCL) frameworks, employing parameter decomposition and various cross-edge strategies, address knowledge retention across devices in multiple federated learning systems, balancing memory and adaptation for new tasks [97].

### v4

@@ II-B Distributed Intelligence Framework
Distributed intelligence frameworks optimize performance by strategically coordinating computational resources across heterogeneous infrastructure, leveraging both proximity-based edge processing and scalable cloud capabilities to address unique constraints and opportunities inherent in collaborative environments. The distributed intelligence framework in ECCC adheres to three key principles specifically adapted for edge-cloud deployment: task parallelism, data parallelism, and model parallelism.
Task parallelism in edge-cloud contexts executes different parts of complex AI tasks concurrently on separate devices, considering edge device limitations and cloud processing capabilities [54], represented as $T=\{t_{1}^{edge},t_{2}^{cloud},...,t_{n}^{hybrid}\}$, where tasks are strategically assigned based on latency requirements, computational complexity, and data locality. Data parallelism trains identical models on different data subsets distributed across multiple edge devices while maintaining privacy [11], expressed as $D=\{d_{1}^{local},d_{2}^{local},...,d_{m}^{local}\}$, where data remains decentralized to address privacy concerns and network bandwidth limitations. Model parallelism deploys different parts of large models on separate devices within the edge-cloud continuum [55], formulated as $M=\{m_{1}^{edge},m_{2}^{fog},...,m_{p}^{cloud}\}$, where model partitioning considers device capabilities, network latency, and energy constraints.
Key design challenges include managing heterogeneous device capabilities across the edge-cloud spectrum, where edge devices have limited computational power but offer low latency, while cloud resources provide extensive computation but introduce communication delays. Network reliability and bandwidth variability require adaptive model partitioning strategies that can dynamically adjust based on real-time network conditions. Energy efficiency becomes paramount for battery-powered edge devices, necessitating intelligent workload distribution that balances computational load with energy consumption. Privacy and data locality requirements demand sophisticated federated learning approaches that enable collaborative intelligence without centralizing sensitive data.
Edge intelligence focuses on real-time processing with sub-10ms latency requirements [56], making it suitable for time-critical applications like autonomous driving and industrial control. Cloud intelligence handles complex computations and large-scale training using virtually unlimited resources [57], enabling sophisticated model development and global knowledge aggregation. Model partitioning strategies must balance workloads by considering model complexity, data size, network topology, and resource constraints specific to edge-cloud deployments [58]. Resource allocation algorithms dynamically assign computational, storage, and communication resources based on real-time demands, device capabilities, and edge-cloud topology [59]. Performance optimization techniques like model compression enhance efficiency across the distributed architecture by reducing communication overhead and enabling deployment on resource-constrained devices [60]. Additionally, fault tolerance mechanisms ensure system reliability despite potential edge device failures, network partitions, or cloud service interruptions, requiring redundancy strategies and graceful degradation protocols.
@@ II-C Resource Types and Constraints
Fundamental system resources, including computation, communication bandwidth, and storage capacity, present inherent constraints that significantly impact distributed architecture design and performance optimization across heterogeneous infrastructure [48]. Specifically, computation resources, denoted as $C_{k}$ for node $k$, represent processing power; communication bandwidth, $B_{ij}$, limits data transfer rates between nodes $i$ and $j$, directly affecting latency; and storage capacity, $S_{k}$, dictates the data volume node $k$ can store, influencing data-intensive applications [61]. Managing these constrained and heterogeneous resources across diverse devices, under dynamic workloads and with limited edge resources, especially energy for battery-powered devices [49], presents significant challenges.
@@ III Model Optimization for Edge-Cloud Systems
Collaborative computing environments present distinctive optimization challenges that fundamentally differ from traditional centralized or purely cloud-based approaches. Distributed environments impose specific constraints that significantly impact optimization strategies: (1) Extreme resource heterogeneity, edge devices range from smartphones with limited processing power to industrial gateways with substantial computational capabilities, necessitating adaptive optimization techniques that can dynamically scale across this spectrum; (2) Intermittent connectivity, network links between edge and cloud are often unreliable, requiring optimization methods that can function during disconnection periods and efficiently synchronize when connectivity is restored; (3) Real-time processing requirements, many edge applications demand ultra-low latency responses (<< 10ms), making traditional batch optimization approaches unsuitable; (4) Privacy-preserving constraints, sensitive data often cannot leave edge devices, requiring optimization techniques that work with limited data visibility; (5) Energy limitations, battery-powered edge devices necessitate energy-aware optimization that balances computational efficiency with power consumption. Consequently, model optimization techniques must be specifically adapted to address these edge-cloud specific constraints, moving beyond general-purpose optimization to develop specialized approaches that leverage the unique characteristics of the edge-cloud continuum while mitigating its inherent limitations.
@@ III-A Model Compression
Resource-constrained environments demand sophisticated model compression techniques to enable deployment of complex AI models while maintaining acceptable performance levels despite limited computational power, memory, and energy resources. Advanced compression approaches [2, 65, 66] have emerged as critical solutions to reduce model size and computational demands while preserving performance. Table II provides a comparison of recent approaches across three primary compression categories: pruning, quantization, and knowledge distillation.
Model compression techniques in ECCC systems can be systematically categorized based on their compression mechanisms and deployment characteristics: (1) Structural modification approaches (pruning) that reduce model complexity by eliminating redundant parameters or architectural components, particularly effective for over-parameterized networks commonly deployed in edge scenarios; (2) Numerical precision optimization (quantization) that maintains model structure while reducing parameter precision requirements, enabling efficient deployment on specialized edge hardware with quantization support; (3) Knowledge transfer and distillation that creates compact models through supervised learning from larger teacher models, providing architectural flexibility and optimal accuracy-efficiency trade-offs for edge deployment. The classification framework reflects both the underlying technical mechanisms and practical deployment considerations essential for edge-cloud environments.
Each compression category exhibits distinct advantages and limitations for edge-cloud deployment scenarios. Pruning techniques excel in reducing model size and computational complexity by eliminating redundant parameters, making them particularly effective for convolutional neural networks where structured pruning can remove entire channels or layers. However, pruning may require fine-tuning to recover accuracy losses and can be sensitive to the pruning ratio selection [67]. Quantization methods provide significant memory reduction and accelerated inference by reducing parameter precision, with minimal accuracy degradation when carefully implemented. They are especially suitable for deployment on specialized hardware with quantization support, though they may face challenges with extremely low-bit quantization or complex model architectures [68]. Knowledge distillation offers the unique advantage of training compact models from scratch while leveraging teacher model knowledge, enabling flexible student model architectures. However, it requires access to both teacher models and training data, making it computationally intensive during the training phase [69].
The choice of compression technique should consider several factors: (1) Resource constraints, pruning is preferred for memory-limited devices, quantization for computational acceleration, and knowledge distillation for balanced performance-efficiency trade-offs; (2) Model complexity, structured pruning works well for over-parameterized models, quantization suits models with redundant precision, and knowledge distillation excels for complex teacher models; (3) Deployment requirements, pruning enables immediate deployment, quantization requires hardware compatibility, and knowledge distillation demands training infrastructure; (4) Performance sensitivity, knowledge distillation typically achieves the best accuracy preservation, followed by quantization and pruning respectively.
Notably, pruning techniques demonstrate significant improvements in inference latency and energy efficiency by strategically eliminating less important parameters [70], as further detailed in Sec. III-A1. Meanwhile, quantization methods, explored in Sec. III-A2, reduce parameter precision requirements, with implementations achieving remarkable energy efficiency in healthcare applications. Knowledge distillation approaches, discussed in Sec. III-A3, enable smaller models to effectively mimic larger counterparts, with recent innovations specifically optimized for efficient LLM inference across edge-cloud boundaries. Across these diverse techniques, a common pattern emerges wherein most successful implementations leverage both edge and cloud resources collaboratively, highlighting the synergistic nature of modern compression strategies in distributed computing environments.
TABLE II: Summary of model compression techniques.
@@ III-A1 Pruning
Parameter elimination strategies reduce model complexity by removing redundant components, enabling efficient AI deployment on resource-constrained devices through targeted removal of less critical weights or architectural elements [71]. We explore various strategies targeting different architectural aspects, such as pruning entire neurons or filters based on performance contribution, as in the two-phase split computing framework (TSCF) [67], or focusing on individual weights. Applied during or after training, dynamic and static pruning balance accuracy, and computational cost. For instance, Hybrid SD [71] employs structural pruning of the Stable Diffusion Model’s U-Net to maintain image quality on edge devices, while the adaptive pruning-split federated learning (PSFed) method [72] optimizes semantic coder updates in a SemCom-assisted SEC framework, reducing training delay and energy consumption.
@@ III-A2 Quantization
Precision reduction techniques effectively minimize memory footprint and computational cost by employing lower-precision representations of model parameters [74], making quantized models particularly suitable for resource-constrained devices. Instead of 32-bit floating-point representations, quantization employs lower precision formats such as 8-bit integers, consequently diminishing memory requirements and accelerating computations on these lower-precision numbers. For example, [73] explored model quantization for edge implementation in arrhythmia monitoring to reduce energy consumption. Additionally, QuAsyncFL [68] integrates quantization with asynchronous federated learning to enhance communication efficiency in cloud-edge-terminal collaborations within the AIoT domain.
@@ III-A3 Knowledge Distillation
Knowledge transfer approaches enable compact ”student” models to achieve comparable performance with reduced complexity by learning from larger ”teacher” models [78], facilitating efficient deployment in resource-constrained environments [65]. Instead of learning from ground truth labels, the student model learns the teacher model’s output distribution, providing richer information than hard labels and enabling more nuanced representations and better generalization [69]. KD benefits ECCC by placing the teacher model in the resource-rich cloud and the distilled student model on edge devices [6], optimizing overall system performance. Agglomerative federated learning (FedAgg) [77] uses a bridge sample based online distillation protocol (BSBODP) for knowledge transfer between nodes across different tiers, enabling larger model training while maintaining privacy and flexibility. Guidance-based knowledge transfer (GKT) [78] employs a larger LLM as a ”teacher” to generate guidance prompts for a smaller ”student” model, enhancing accuracy and speed without fine-tuning. Growth-adaptive distillation [6] automatically adjusts the student model structure for efficient, low-power edge deployment. In cloud-edge collaborative reinforcement learning, algorithm-independent knowledge distillation [79] accelerates online RL agent convergence by leveraging pre-trained cloud models and offline agent interactions. A federated bi-directional knowledge distillation strategy [80] enhances fault detection in manufacturing by integrating quality knowledge from both cloud and edge, strengthening interaction between sub-processes and hierarchical levels.
@@ III-B Model Adaptation
Adaptive learning paradigms enable AI models to effectively respond to new tasks, datasets, or evolving data distributions within dynamic distributed computing environments. To achieve this adaptability, three primary learning paradigms are widely employed: transfer learning (TL), which accelerates learning on target tasks by leveraging knowledge from pre-trained models; federated learning (FL), which facilitates privacy-preserving, distributed training across multiple clients; and continual learning (CL), which enables models to learn sequentially from evolving data streams without catastrophic forgetting. The interplay and distinct roles of these paradigms in fostering adaptability within ECCC systems are conceptually depicted in Fig. 4.
While some taxonomies treat federated learning as a distinct resource management or privacy technique, our classification positions FL within model adaptation because it fundamentally addresses how models adapt to distributed data patterns and heterogeneous environments, which is a core adaptation challenge in ECCC systems. Specifically: (1) Domain adaptation techniques (transfer learning) enable models to adapt across different domains or tasks by leveraging pre-existing knowledge; (2) Distributed adaptation methods (federated learning) enable models to adapt collaboratively across multiple data sources while preserving privacy and handling non-IID distributions; (3) Temporal adaptation approaches (continual learning) enable models to adapt incrementally to evolving data and new tasks over time. The classification emphasizes the adaptation aspect of each technique while acknowledging that FL also serves privacy and resource distribution functions, making it inherently multi-faceted in ECCC systems.
Each learning approach offers unique advantages for different edge-cloud deployment scenarios and data distribution characteristics. Transfer learning excels in scenarios with limited target domain data by leveraging pre-trained models from related domains, achieving faster convergence and reduced computational requirements during training [81]. However, transfer learning performance heavily depends on the similarity between source and target domains, potentially struggling with domain shifts or highly specialized tasks. Federated learning provides exceptional privacy preservation by keeping data distributed across edge devices while enabling collaborative model training, making it ideal for scenarios with sensitive data or regulatory constraints [77]. Nevertheless, FL faces challenges with non-IID data distributions, communication overhead, and convergence difficulties in heterogeneous environments [82]. Continual learning offers superior adaptability to evolving data distributions and new tasks without requiring access to historical data, enabling lifelong learning capabilities crucial for dynamic edge environments [83]. However, CL techniques often struggle with catastrophic forgetting and require sophisticated regularization mechanisms to maintain performance on previous tasks.
The choice of learning paradigm should align with specific data distribution characteristics: (1) Homogeneous data distributions, transfer learning works well when source and target domains share similar characteristics, while federated learning benefits from consistent data patterns across participants; (2) Non-IID data distributions, specialized FL techniques (e.g., FedProx, FedNova) address statistical heterogeneity, while transfer learning may require domain adaptation methods; (3) Temporal data shifts, continual learning excels in adapting to evolving data patterns, while transfer learning may need periodic retraining; (4) Limited local data, transfer learning provides quick adaptation with few samples, while FL can leverage collective knowledge across devices even with sparse local datasets. The optimal selection often involves hybrid approaches that combine multiple paradigms to address specific edge-cloud requirements and constraints. Furthermore, Table III presents a comparative summary of recent research efforts utilizing TL, FL, and CL, highlighting their specific approaches, application domains, and system characteristics within ECCC. Following this, the subsequent sections delve into the transfer learning in Sec. III-B1, federated learning in Sec. III-B2, and continual learning in Sec. III-B3.
@@
Method Year Venue Type Domain Edge Cloud ECC MC
@@
Note for ’Type’: TL = Transfer Learning, FL = Federated Learning, CL = Continual Learning, MC = Model Compression.
@@ III-B1 Transfer Learning
Pre-trained model adaptation to new tasks or datasets leverages transfer learning principles to reduce training time and data requirements, thereby improving deployment efficiency in resource-constrained environments [81]. For example, in surface defect detection, a model pre-trained on a general image dataset can be fine-tuned with a smaller, specialized dataset to identify specific defects [81]. Similarly, in cloud-device collaborative learning for multimodal large language models (MLLMs), techniques like adapter-based knowledge distillation (AKD) facilitate knowledge transfer from larger cloud models to smaller, compressed device models [69], maintaining performance while minimizing computational burden. In cloud-edge collaborative fault diagnosis, transfer learning addresses data heterogeneity and distribution shifts by fine-tuning source models trained on diverse user data to bridge distribution gaps and enable effective diagnosis on target datasets [87].
III-B2 Model Adaptation in Federated Learning
The explosive growth of FL research has made it an indispensable paradigm in distributed AI systems, driving critical need to explore how model adaptation mechanisms can be effectively implemented within federated frameworks [100]. As one of the most active research domains in distributed machine learning, FL cannot be overlooked when examining model adaptation techniques, particularly given its unique challenges in handling heterogeneous data distributions and diverse edge environments [92]. We focus specifically on model adaptation strategies within FL contexts, exploring how adaptive mechanisms enable models to effectively respond to distributed, non-IID data patterns while maintaining collaborative training capabilities [94]. Recent frameworks demonstrate sophisticated adaptation approaches, such as FedCAE [98] for machine fault diagnosis that adapts models to diverse industrial environments, and QuAsyncFL [68] for AIoT applications that enables adaptive synchronization strategies across heterogeneous devices. Advanced adaptation mechanisms like SFETEC [106] optimize model adaptation for thing-edge-cloud environments by implementing adaptive model splitting strategies that dynamically adjust to varying communication constraints and device capabilities, addressing non-IID data challenges through intelligent partitioning. Security-aware adaptation frameworks such as FlexibleFL [99] integrate adaptive evaluation mechanisms that assess uploaded parameters and participant contributions, enabling models to adapt against evolving poisoning attacks while maintaining collaborative learning effectiveness. Furthermore, adaptive participant selection and topology optimization strategies [96] enhance FL efficiency by dynamically adjusting collaboration patterns, while blockchain integration approaches [82] provide adaptive security mechanisms that evolve with emerging threats, demonstrating how model adaptation principles are fundamental to advancing FL systems in distributed edge-cloud environments.
@@ III-B3 Continual Learning
Incremental adaptation strategies enable models to evolve with dynamic data and emerging tasks while addressing catastrophic forgetting challenges through lifelong learning capabilities [83]. Several strategies, including regularization, replay, and dynamic architectures, mitigate this issue, ensuring robust and adaptable systems. Specifically, continual learning enables digital twins to synchronize with their evolving physical counterparts and sensor data in edge computing, providing accurate predictions [103, 102]. For instance, in intelligent transportation systems (ITS), cloud-edge collaborative continual adaptation frameworks, leveraging visual prompts and knowledge distillation, enhance the generalization of lightweight edge models for object detection in dynamic environments [104]. Similarly, online management frameworks, such as EdgeC3, optimize model aggregation and dynamic offloading for continuous data streams in edge-cloud collaborative continual learning [105]. Cross-edge federated continual learning (Cross-FCL) frameworks, employing parameter decomposition and various cross-edge strategies, address knowledge retention across devices in multiple FL systems, balancing memory and adaptation for new tasks [107].
III-B4 Split Learning
Distributed learning paradigms partition neural network models across multiple computational nodes to enable collaborative training without sharing raw data, addressing computational limitations while preserving privacy [108]. Unlike FL where complete models are trained locally, split learning divides the neural network architecture at specific cut layers, with edge devices typically handling initial feature extraction layers while cloud or edge servers process deeper computational layers [109]. Strategic architectural division effectively addresses the computational limitations of resource-constrained IoT devices while preserving data privacy, as demonstrated in eHealth applications where heterogeneous workload-based resource recommendation models leverage federated split learning (FSL) for smart city deployments [110]. Advanced frameworks like MADRL-based model partitioning systems optimize the collaboration between cloud-edge-device architectures by jointly considering model aggregation control and resource allocation, addressing device heterogeneity through adaptive partitioning strategies rather than uniform approaches [111].
The optimal selection of split points significantly impacts system performance across multiple dimensions including computational load distribution, communication overhead, and privacy preservation. Efficient parallel split learning (EPSL) frameworks demonstrate how intelligent cut layer selection, combined with subchannel allocation and power control optimization, can minimize per-round training latency in resource-constrained wireless edge networks [112]. The SFETEC framework exemplifies optimized split learning for thing-edge-cloud environments, where base models are trained at client devices while core models are processed at edge servers and aggregated at the cloud, significantly reducing communication overhead and training duration compared to traditional federated approaches [106]. Advanced split learning architectures like Split-and-Bridge address class incremental learning challenges by adaptively partitioning neural networks into specialized segments for handling new and old tasks, demonstrating superior performance in knowledge distillation-based continual learning scenarios [113].
Split learning implementations must carefully balance communication efficiency with model performance, particularly in resource-constrained environments. The CSE-FSL framework introduces auxiliary networks for local weight updates while maintaining single server-side models, dramatically reducing gradient transmission requirements and server storage overhead [114]. For wearable IoMT devices, FedSL frameworks enable collaborative healthcare analytics by allowing even low-end devices to participate in model training through strategic model partitioning, while maintaining privacy by avoiding raw data transmission over wireless networks [115]. Comprehensive evaluations comparing FL and split learning in IoT settings reveal that while split learning provides better learning performance under imbalanced data distributions, FL demonstrates superior communication efficiency, highlighting the importance of application-specific optimization strategies [108]. Optimization considerations collectively enable split learning deployment across diverse 6G edge networks, facilitating connected intelligence through distributed model training architectures [109].